OpenStudy (sonyalee77):

The ratio of Brenda’s age to Jason’s age is 4:5. In eight years the ratio will be 6:7. How old is Brenda now?

OpenStudy (tkhunny):

Step #1 - Name things. B is Brenda's Age Now J is Jason's Age Now Okay, now we can talk about them. If their current ages have a ratio of 4:5, can you express that in mathematical notation? Hint: Think Ratio = Fraction.

OpenStudy (kropot72):

J=5/4 * B = 1.25B .............(1) \[\frac{B+8}{J+8}=\frac{6}{7}\ ............(2)\] Cross multiplying (2) gives 7B + 56 = 6J + 48 ..................(3) Plugging the value for J from (1) into (3) we get 7B + 56 = 7.5B + 48 ..............(4) Now you just need to solve (4) to find B.

OpenStudy (kropot72):

@Sonyalee77 Do you understand?

OpenStudy (sonyalee77):

no

OpenStudy (sonyalee77):

@campbell_st

OpenStudy (kropot72):

7B + 56 = 7.5B + 48 ..............(4) Now you just need to solve (4) to find B. First subtract 7B from both sides to get 7.5B - 7B + 48 = 56 which simplifies to 0.5B + 48 = 56 Now subtract 48 from both sides to get 0.5B = 8 ............(5) Now if you divide both sides of equation (5) by 0.5 you will get B, Brenda's present age.

OpenStudy (kropot72):

@Sonyalee77 Are you there?

OpenStudy (sonyalee77):

is 16 right?

OpenStudy (kropot72):

Yes, Brenda's age now is 16 years.